# Long-Context Language Modeling with Parallel Context Encoding

### v1

@@ 7 Ablation Studies
We conduct comprehensive ablations to show the effectiveness of our training data mixture, pre-training and fine-tuning the encoder, and the warmup training stage. We also ablate to verify the effectiveness of the KL divergence loss in CEPED.
@@ 7.1 Training Settings
Even though CEPE was trained with long documents, it also achieves strong performance on retrieval-augmented applications. In this subsection, we study a different data strategy: training CEPE with retrieved documents. Specifically, we pair the training sequences described in §2.2 with retrieved passages from the retrieval split of RP using Contriever. We use the first 256 tokens of the decoder input $X$ as the query and retrieve $k=16$ passages to form the additional contexts $\mathcal{C}$. Then, we train CEPE with retrieved passages (RetDoc) using the same training settings as in §2.3.
As shown in Table 5, we find that training on RetDoc results in slightly stronger performances in the retrieval-augmentation setting, but the results on long documents are worse. Augmenting a pre-training corpus with retrieval contexts can be extremely computationally and storage expensive. CEPE’s simple long-document data strategy achieves a good balance between efficient training and strong performance on both long-context and retrieval-augmented applications.
@@
Encoder training.
Finally, we investigate how to best train the encoder. To this end, we train CEPE with (1) freezing the encoder after pre-training and the warmup stage, (2) training with a randomly initialized encoder, and (3) using the pre-trained model without the warmup stage. As shown in Table 5, we find that the copying warmup stage and fine-tuning the encoder during training are both crucial for strong performance.
@@ 7.2 KL Divergence
The key component of CEPED is the KL Divergence loss. To understand the importance of this auxiliary loss, we explore the performance of CEPED when trained without the KL Divergence loss as well as with difference coefficients for each loss. Let $\mathcal{L}_{CE},\mathcal{L}_{KL}$ be the cross entropy loss and the KL Divergence loss, respectively. Then, the total loss is $\mathcal{L}=c_{CE}\mathcal{L}_{CE}+c_{KL}\mathcal{L}_{KL}$. The results are shown in Table 13. We find that the KL Divergence loss is especially crucial for summarization tasks and QALT.
@@ 8 Related Work
Many recent works on long-context language models aim to solve the problem of positional embedding extrapolation in transformers Peng et al. (2023); Chen et al. (2023a). Others fine-tune LMs on longer sequences Xiong et al. (2023); Chen et al. (2023b); Together (2023a) or compress the context into shorter forms Choromanski et al. (2021); Chevalier et al. (2023). Notably, several recent papers propose to extend the context window of LMs by modifying the attention mechanism: Xiao et al. (2023) discover the use of “sink tokens” in sliding windows and Bertsch et al. (2023) retrieve relevant tokens from a cache instead of attending to all tokens. This results in memory-efficient long-context LMs, but they have diminishing returns with longer contexts, as the same positional embedding may be seen multiple times and they can not fully utilize all tokens. The key advantage of CEPE is that it does not degrade for inputs longer than the training length while achieving great efficiency compared to full fine-tuning approaches.
Novel architectures and pre-training methods, such as S4 Gu et al. (2022), RPT Rubin and Berant (2023), and Mamba Gu and Dao (2023), also extend the context window at greater efficiency. However, pre-training is extremely expensive at scale and these methods cannot leverage existing powerful pre-trained LLMs. It is also unclear if state-space models can achieve comparable performance with transformers Jelassi et al. (2024).
@@
Augmenting LMs with retrieval has been useful in many applications, such as open-domain question answering. Recently, combining LMs with retrieval systems for more generalized purposes, such as language modeling, has been explored: Guu et al. (2020); Borgeaud et al. (2022); Izacard et al. (2022b); Min et al. (2023) pre-train LMs with retrieval, and Shi et al. (2023b); Lin et al. (2023) use logits interpolation from separate forward passes to incorporate retrieval information.
@@ Acknowledgements
We want to acknowledge Sadhika Malladi, Alexander Wettig, Mengzhou Xia, Dan Friedman, Adithya Bhaskar, Victoria Graf, and other members of the Princeton NLP group for their useful feedback and discussion. Tianyu Gao is supported by an IBM PhD Fellowship. This work is also gratefully supported by an NSF CAREER Award (IIS-2239290) and Intel.
@@ A.2 Training CEPE
The attention module in LLaMA-2 consists of four projection matrices: key, value, query, and output. In contrast to original transformers Vaswani et al. (2017), the output projection matrix is used as an additional attention output projection. When we first insert the cross-attention layers into the decoder, we initialize the weights of the key, value, and query projection matrices with the respective weights from the decoder’s self-attention layer in the same transformer block. Furthermore, since the hidden dimension of the encoder is smaller than the hidden dimension of the decoder, $d<D$, we use only copy the first $d$ rows of the key and value projection matrices from the self-attention module to the cross-attention module. Lastly, the output projection matrix is initialized with all zeros. While we did not investigate this initialization in detail, the intuition is that we want the model should use the tokens from the encoder using a similar mechanism as the decoder uses for its own tokens. However, we want the model to learn the output projection from scratch, as it may be too disruptive to have doubled the number of attention modules.
@@ A.3 Training CEPED
We leverage an additional distillation loss for CEPED. For encoder input $C$ and decoder input $X$, we first calculate the logits $\theta(\text{concat}(C,X))$ by running forward passes with the original model parameterized by $\theta$. Due to storage constraints, we only store the top 50 likelihoods and their indices in the vocabulary for each token in $X$, following Askell et al. (2021); Bai et al. (2022).
@@
[…]
@@ Appendix B Baseline Implementations
Although RePlug Shi et al. (2023b) was introduced as a method to augment language models with retrieval, we found that the technique of interpolating logits from separate forward passes can also transfer well to the long context setting. Among the methods that we compare to, RePlug uniquely improves performance upon the base model in both the long-context and the retrieval-augmented LM settings. This gives us an additional point of comparison across the two settings.
@@
We follow the implementation of StreamingLLM from the original authors [^6] https://github.com/mit-han-lab/streaming-llm Xiao et al. (2023). Specifically, we use their best settings, where we enable the positional shifts and cache 4 sink tokens and 2044 recent tokens.
@@ C.2 In-context Learning
For our in-context learning experiments, we use the datasets commonly used in previous works Zhao et al. (2021); Lu et al. (2022); Han et al. (2023); Ratner et al. (2023): SST-2 Socher et al. (2013), MovieReview (MR Pang and Lee, 2005), AGNews Zhang et al. (2015), SST-5 Socher et al. (2013), TREC Voorhees and Tice (2000), DBPedia Zhang et al. (2015), NLU Liu et al. (2019a), BANKING77 Casanueva et al. (2020), CLINIC150 Larson et al. (2019). We follow the prompts used in Ratner et al. (2023) for all datasets. During evaluation, we first calculate the log-likelihood of each option and select the option with the highest likelihood. We sample the in-context learning demonstrations from the training set such that each label has an equal number of demonstrations (except for possible reminders).
@@ C.3 ZeroScrolls
Appendix D Ablations
D.1 Training Settings
We show the full results for the ablation studies in Table 11 and 12, and 13.

### v2

@@ 7 Ablation Studies
We conduct comprehensive ablations to show the effectiveness of our training data mixture, pre-training and fine-tuning the encoder, and the warmup training stage. We also ablate to verify the effectiveness of the KL divergence loss in CEPED. Lastly, we evaluate on Needle in the Haystack Kamradt (2024), and analyze the results in §D.
@@ 7.1 Training Settings
Even though CEPE was trained with long documents, it achieves strong performance on retrieval-augmented applications. We also test a different data strategy: training CEPE with retrieved documents using the training settings from §2.3.
As shown in Table 5, we find that training on retrieved documents (RetDoc) results in slightly stronger performances in the retrieval-augmentation setting, but performs worse on long documents. Augmenting a pre-training corpus with retrieval contexts is extremely computationally and storage expensive. CEPE’s simple long-document data strategy achieves a good balance between efficient training and strong performance on both long-context and retrieval-augmented applications.
@@ 7.2 KL Divergence
The key component of CEPED is the KL Divergence loss. To understand the importance of this auxiliary loss, we explore the performance of CEPED when trained without the KL Divergence loss as well as with difference coefficients for each loss. Results are shown in Table 14, and we find that the KL Divergence loss is crucial for summarization tasks and QALT. More details are in §E.2.
@@ 8 Related Work
Many recent works on long-context LMs aim to solve the problem of positional embedding extrapolation in transformers Peng et al. (2024); Chen et al. (2023). Others fine-tune LMs on longer sequences Xiong et al. (2023); Chen et al. (2024); Together (2023a) or compress the context into shorter forms Yoshida et al. (2020); Choromanski et al. (2021); Chevalier et al. (2023). Notably, several recent papers propose to extend the context window of LMs by modifying the attention mechanism: Xiao et al. (2024b) discover the use of “sink tokens” in sliding windows, and Bertsch et al. (2023); Xiao et al. (2024a) retrieve relevant tokens from a cache instead of attending to all tokens. This results in memory-efficient long-context LMs, but they have diminishing returns with longer contexts, as the same positional embedding may be seen multiple times and they can not fully utilize all tokens. The key advantage of CEPE is that it does not degrade for inputs longer than the training length while achieving great efficiency compared to full fine-tuning approaches. Techniques have been designed for specific applications, such as for in-context learning Ratner et al. (2023); Hao et al. (2022), but we focus on general long-context language modeling.
Novel architectures and pre-training methods, such as S4 Gu et al. (2022), RPT Rubin and Berant (2023), YOCO Sun et al. (2024), and Mamba Gu and Dao (2023), also extend the context window at greater efficiency. However, pre-training is extremely expensive at scale and these methods cannot leverage existing powerful pre-trained LLMs. It is also unclear if state-space models can achieve comparable performance with transformers Jelassi et al. (2024).
@@
Augmenting LMs with retrieval has been useful in many applications, such as open-domain question answering. Recently, combining LMs with retrieval systems for more generalized purposes has been explored: Guu et al. (2020); Borgeaud et al. (2022); Izacard et al. (2022b); Min et al. (2023) pre-train LMs with retrieval, and Shi et al. (2024); Lin et al. (2024) use logits interpolation from separate forward passes to incorporate retrieval information.
@@ Acknowledgements
We want to acknowledge Sadhika Malladi, Alexander Wettig, Mengzhou Xia, Dan Friedman, Amanda Bertsch, Adithya Bhaskar, Victoria Graf, and other members of the Princeton NLP group for their useful feedback and discussion. Tianyu Gao is supported by an IBM PhD Fellowship. This work is also gratefully supported by an NSF CAREER Award (IIS-2239290) and Intel.
@@ A.2 Training CEPE
The attention module in LLaMA-2 consists of four projection matrices: key, value, query, and output. In contrast to original transformers Vaswani et al. (2017), the output projection matrix is used as an additional attention output projection. When we first insert the cross-attention layers into the decoder, we initialize the weights of the key, value, and query projection matrices with the respective weights from the decoder’s self-attention layer in the same transformer block. Furthermore, since the hidden dimension of the encoder is smaller than the hidden dimension of the decoder, $d<D$, we use only copy the first $d$ rows of the key and value projection matrices from the self-attention module to the cross-attention module. Lastly, the output projection matrix is initialized with all zeros. While we did not investigate this initialization in detail, the intuition is that we want the model to use the tokens from the encoder using a similar mechanism as the decoder uses for its own tokens. However, we want the model to learn the output projection from scratch, as it may be too disruptive to have doubled the number of attention modules.
@@ A.3 Training CEPED
We leverage an additional distillation loss for CEPED. For encoder input $\mathcal{C}$ and decoder input $X$, we first calculate the logits $\mathcal{M}_{\textrm{dec}}(\text{concat}(\mathcal{C},X))$ by running forward passes with the original model parameterized by $\mathcal{M}_{\textrm{dec}}$. Due to storage constraints, we only store the top 50 likelihoods and their indices in the vocabulary for each token in $X$, following Askell et al. (2021); Bai et al. (2022).
@@
$\mathcal{L}_{\textrm{KL}}$ $=D_{\textrm{KL}}(\mathcal{M}_{\textrm{dec}}(S)||\mathcal{M}_{\mathrm{CEPE}}(\mathcal{C},X))$
where $S=\text{concat}(\mathcal{C},X)$, $\mathcal{M}_{\textrm{dec}}(S)$ is the probability distribution of the top 50 tokens for $X$, and $\mathcal{M}_{\mathrm{CEPE}}(\mathcal{C},X)$ takes $\mathcal{C}$ as the encoder input and $X$ as the decoder input and outputs the probability distribution of the same 50 tokens on $X$.
@@ Appendix B Baseline Implementations
Although RePlug Shi et al. (2024) was introduced as a method to augment language models with retrieval, we found that the technique of interpolating logits from separate forward passes can also transfer well to the long context setting. Among the methods that we compare to, RePlug uniquely improves performance upon the base model in both the long-context and the retrieval-augmented LM settings. This gives us an additional point of comparison across the two settings.
@@
We follow the implementation of StreamingLLM from the original authors [^7] https://github.com/mit-han-lab/streaming-llm Xiao et al. (2024b). Specifically, we use their best settings, where we enable the positional shifts and cache 4 sink tokens and 2044 recent tokens.
@@ C.2 In-context Learning
For our in-context learning experiments, we use the datasets commonly used in previous works Zhao et al. (2021); Lu et al. (2022); Han et al. (2023); Ratner et al. (2023): SST-2 Socher et al. (2013), MovieReview (MR Pang and Lee, 2005), AGNews Zhang et al. (2015), SST-5 Socher et al. (2013), TREC Voorhees and Tice (2000), DBPedia Zhang et al. (2015), NLU Liu et al. (2019a), BANKING77 Casanueva et al. (2020), CLINIC150 Larson et al. (2019). We follow the prompts used in Ratner et al. (2023) for all datasets. During evaluation, we first calculate the log-likelihood of each option and select the option with the highest likelihood. We sample the in-context learning demonstrations from the training set such that each label has an equal number of demonstrations (except for possible remainders).
@@ C.3 ZeroScrolls
Dataset Mean Median Max. Min. $l\in[4K,32K]$ (%) $l>32K$ (%)
NarrativeQA 75998.7 69163 264143 18260 5.8 94.2
QASPER 5278.5 4896.0 17626 2003 61.6 0.0
QuALITY 8035.8 8285.5 10779 3443 96.2 0.0
GovReport 11345.4 10128.5 48074 2441 91.2 0.6
SummScreenFD 9924.2 9364.0 27565 3078 99.1 0.0
QMSum 16027.7 14811.5 34543 3120 94.1 4.8
Table 11: ZeroScrolls length statistics. $l$ is the number of input tokens in each example. We report the mean, median, maximum, and minimum number of input tokens in each dataset. We also report the percentage of examples that have between $4,096$ and $32,768$ tokens ($l\in[4K,32K]$) and the percentage of examples that have over $32,768$ tokens ($l>32K$). We observe that only NarrativeQA has a substantial number of test examples with more than 32K tokens, making it the most useful for evaluating long-context language models.
Appendix D Needle in the Haystack
Figure 4: Needle in the Haystack evaluation on CEPE-LLaMA-2. The white dotted line denotes the training length.
Needle in the Haystack Kamradt (2024) is a synthetic task designed to test in-context retrieval abilities of long-context language models. In this task, a needle (e.g., “The best thing to do in San Francisco is to eat a sandwich and sit in Dolores Park on a sunny day”) is placed in a long text of essays, and the model is tasked with retrieving the needle from the context—that is, the model is expected to answer “What is the best thing to do in San Francisco?” with the needle.
We evaluate CEPE-LLaMA-2, where $2K$ tokens are used in the decoder, and the rest of the context is used in the encoder. For this experiment, we also scale the cross-attention scores with the ratio between the number of tokens during inference and the number of tokens during training, following Peng et al. (2024). We follow the evaluation setting of Fu et al. (2024), but split the context by sentence boundaries when inputting additional tokens to the encoder. We find that while CEPE-LLaMA-2 is able to perfectly retrieve the needle in context at the training sequence length (~10K tokens) and at some longer sequences (~14K tokens), it struggles at other lengths. This is likely due to the training/inference discrepancies on the input length and we will explore training objectives that can mitigate such discrepancies in the future. Other efficient methods, such as StreamingLLM Xiao et al. (2024b), also struggle with copying tokens that are far from the current context.
Appendix E Ablations
We show the full results for the ablation studies in Table 12 and 13, and 14. The subsections below describe the ablation settings in more detail.
E.1 Training Settings
Training with retrieved documents.
In this subsection, we study training CEPE with retrieved documents. Specifically, we pair the training sequences described in §2.2 with retrieved passages from the retrieval split of RP using Contriever. We use the first 256 tokens of the decoder input $X$ as the query and retrieve $k=16$ passages to form the additional contexts $\mathcal{C}$. Then, we train CEPE with retrieved passages (RetDoc) using the same training settings as in §2.3.
Encoder training.
We also investigate how to best train the encoder. To this end, we train CEPE with (1) freezing the encoder after pre-training and the warmup stage, (2) training with a randomly initialized encoder, and (3) using the pre-trained model without the warmup stage. As shown in Table 5, we find that the copying warmup stage and fine-tuning the encoder during training are both crucial for strong performance.
E.2 KL Divergence
The key component of CEPED is the KL Divergence loss. To understand the importance of this auxiliary loss, we explore the performance of CEPED with changes to the loss function in this subsection. Let $\mathcal{L}_{CE},\mathcal{L}_{KL}$ be the cross entropy loss and the KL Divergence loss, respectively. Then, the total loss is $\mathcal{L}=c_{CE}\mathcal{L}_{CE}+c_{KL}\mathcal{L}_{KL}$, where $c_{CE}$ and $c_{KL}$ are coefficients for the cross entropy loss and the KL Divergence loss, respectively. We vary the coefficients $c_{CE}$ and $c_{KL}$ to study the importance of the KL Divergence loss. The results are presented in Table 14. Without the KL Divergence loss, CEPED may still perform well on NarrativeQA and Qasper, but the performance on the summarization tasks and QuALITY may decrease as a result.